OpenStudy (anonymous):

I need help urgently!! Find zw and z/w. Leave your answers in polar form. z=5-5i w=sqrt(3)+i

OpenStudy (fibonaccichick666):

Do you know what the polar form of z is?

OpenStudy (anonymous):

No.

OpenStudy (fibonaccichick666):

ok, so do you know what polar form looks like?

OpenStudy (anonymous):

Yes, (r,theta).

OpenStudy (fibonaccichick666):

ok, so in polar in terms of complex numbers means it looks like this :\[re^{i\theta}=r(cos\theta+isin\theta)\]

OpenStudy (anonymous):

Okay. How do we begin converting it? Because, I don't understand that part, I mean. \[(5-5i)/(\sqrt{3}+i)\]

OpenStudy (anonymous):

If \[ z=x+yi \]Then... \[ r=\sqrt{x^2+y^2}\\ \theta = \tan^{-1}\left(\frac yx\right) \]

OpenStudy (anonymous):

It's the same \((x,y)\to (r,\theta)\) you learn in Calc

OpenStudy (fibonaccichick666):

thank wio... I was just getting to that point

OpenStudy (anonymous):

which is why I think complex analysis is so much bs.

OpenStudy (fibonaccichick666):

So, if you would graph your point in the complex plane, it would look like this:

OpenStudy (anonymous):

\[r=7.07, \Theta=\frac{ -5 }{ 5 }\] Is that correct?

OpenStudy (fibonaccichick666):

so then you need to find your r, thhis is how we derive it all 

OpenStudy (fibonaccichick666):

so can you find the hypotenuse of the triangle I've drawn?

OpenStudy (anonymous):

Yes, I already did. Sqrt{50}

OpenStudy (fibonaccichick666):

yes, ok so now, what is the exact value of theta?

OpenStudy (anonymous):

-5/5